### v2

@@ 3 Proof of the theorem
BakerM.JensenD.Degeneration of linear series from the tropical point of view and applicationsM. Baker, S.Payne (eds) Nonarchimedean and Tropical GeometrySimons Symposia2016365–433@article{ref1, author = {M. Baker}, author = {D. Jensen}, title = {Degeneration of linear series from the tropical point of view and applications}, journal = {M. Baker, S.Payne (eds) Nonarchimedean and Tropical Geometry}, series = {Simons Symposia}, year = {2016}, pages = {365-433}} Cook-PowellK.JensenD.Tropical methods in hurwitz-brill-noether theoryAdvances in Math.3982022108199@article{ref5, author = {K. Cook-Powell}, author = {D. Jensen}, title = {Tropical methods in Hurwitz-Brill-Noether theory}, journal = {Advances in Math.}, volume = {398}, year = {2022}, pages = {108199}} CoolsF.DraismaJ.PayneS.RobevaE.A tropical proof of the brill-noether theoremAdvances in Math.2302012759–776@article{ref4, author = {F. Cools}, author = {J. Draisma}, author = {S. Payne}, author = {E. Robeva}, title = {A tropical proof of the Brill-Noether Theorem}, journal = {Advances in Math.}, volume = {230}, year = {2012}, pages = {759-776}} CoolsF.PanizzutM.The gonality sequence of complete graphsElectr. J. Comb.24(4)2017P4.1@article{ref6, author = {F. Cools}, author = {M. Panizzut}, title = {The gonality sequence of complete graphs}, journal = {Electr. J. Comb.}, volume = {24(4)}, year = {2017}, pages = {P4.1}} CoppensM.MartensG.Secant spaces and clifford’s theoremComposition Math.781991193–212@article{ref2, author = {M. Coppens}, author = {G. Martens}, title = {Secant spaces and Clifford's Theorem}, journal = {Composition Math.}, volume = {78}, year = {1991}, pages = {193-212}} DeveauA.JensenD.KainicJ.MitropolskyD.Gonality of random graphsInvolve92016715–720@article{ref3, author = {A. Deveau}, author = {D. Jensen}, author = {J. Kainic}, author = {D. Mitropolsky}, title = {Gonality of random graphs}, journal = {Involve}, volume = {9}, year = {2016}, pages = {715-720}} PfluegerN.Special divisors on marked chains of cyclesJ. Combin. Theory Ser. A1502017182–207@article{ref7, author = {N. Pflueger}, title = {Special divisors on marked chains of cycles}, journal = {J. Combin. Theory Ser. A}, volume = {150}, year = {2017}, pages = {182-207}}

### v3

@@ 3 Proof of the theorem
BakerM.JensenD.Degeneration of linear series from the tropical point of view and applicationsNonarchimedean and Tropical Geometry, pp. 365-433. Simons Symp., Springer, Cham (2016)@article{ref1, author = {M. Baker}, author = {D. Jensen}, title = {Degeneration of linear series from the tropical point of view and applications}, journal = {Nonarchimedean and Tropical Geometry, pp. 365-433. Simons Symp., Springer, % Cham (2016)}} Cook-PowellK.JensenD.Tropical methods in hurwitz-brill-noether theoryAdvances in Math.3982022Paper No. 108199, 42 pp.@article{ref5, author = {K. Cook-Powell}, author = {D. Jensen}, title = {Tropical methods in Hurwitz-Brill-Noether theory}, journal = {Advances in Math.}, volume = {398}, year = {2022}, pages = { Paper No. 108199, 42 pp.}} CoolsF.DraismaJ.PayneS.RobevaE.A tropical proof of the brill-noether theoremAdvances in Math.2302012759–776@article{ref4, author = {F. Cools}, author = {J. Draisma}, author = {S. Payne}, author = {E. Robeva}, title = {A tropical proof of the Brill-Noether Theorem}, journal = {Advances in Math.}, volume = {230}, year = {2012}, pages = {759-776}} CoolsF.PanizzutM.The gonality sequence of complete graphsElectr. J. Comb.24(4)2017Paper No. 4.1, 20 pp@article{ref6, author = {F. Cools}, author = {M. Panizzut}, title = {The gonality sequence of complete graphs}, journal = {Electr. J. Comb.}, volume = {24(4)}, year = {2017}, pages = {Paper No. 4.1, 20 pp}} CoppensM.MartensG.Secant spaces and clifford’s theoremCompositio Math.781991193–212@article{ref2, author = {M. Coppens}, author = {G. Martens}, title = {Secant spaces and Clifford's Theorem}, journal = {Compositio Math.}, volume = {78}, year = {1991}, pages = {193-212}} DeveauA.JensenD.KainicJ.MitropolskyD.Gonality of random graphsInvolve92016715–720@article{ref3, author = {A. Deveau}, author = {D. Jensen}, author = {J. Kainic}, author = {D. Mitropolsky}, title = {Gonality of random graphs}, journal = {Involve}, volume = {9}, year = {2016}, pages = {715-720}} PfluegerN.Special divisors on marked chains of cyclesJ. Combin. Theory Ser. A1502017182–207@article{ref7, author = {N. Pflueger}, title = {Special divisors on marked chains of cycles}, journal = {J. Combin. Theory Ser. A}, volume = {150}, year = {2017}, pages = {182-207}}